OpenStudy (anonymous):

-6 + t =-14 t= -8 how do I show my work

OpenStudy (matt101):

Just show things step by step - for this question it's pretty simple: \[-6+t=-14\]\[-6+t+6=-14+6\]\[t=-8\]

OpenStudy (anonymous):

Look in all of equations we have this : there are some fundamental row operations 1) you can interchange rows 2) you can multiply or divide the row with real no 3) you can add subtract multiple of any row in another row . Got it ?:)

OpenStudy (anonymous):

yeah I got it thanks guys

OpenStudy (anonymous):

Realy ????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

shouldn't I say thank you

OpenStudy (anonymous):

wat????

OpenStudy (anonymous):

i should say thank you for helping me

OpenStudy (anonymous):

heh ! You re welcome :)